# DeLiVR: Differential Spatiotemporal Lie Bias for Efficient Video Deraining

### v1

@@ Abstract
Videos captured in the wild often suffer from rain streaks, blur, and noise. In addition, even slight changes in camera pose can amplify cross-frame mismatches and temporal artifacts. Existing methods rely on optical flow or heuristic alignment, which are computationally expensive and less robust. To address these challenges, Lie groups provide a principled way to represent continuous geometric transformations, making them well-suited for enforcing spatial and temporal consistency in video modeling. Building on this insight, we propose DeLiVR, an efficient video deraining method that injects spatiotemporal Lie-group differential biases directly into attention scores of the network. Specifically, the method introduces two complementary components. First, a rotation-bounded Lie relative bias predicts the in-plane angle of each frame using a compact prediction module, where normalized coordinates are rotated and compared with base coordinates to achieve geometry-consistent alignment before feature aggregation. Second, a differential group displacement computes angular differences between adjacent frames to estimate a velocity. This bias computation combines temporal decay and attention masks to focus on inter-frame relationships while precisely matching the direction of rain streaks. Extensive experimental results demonstrate the effectiveness of our method on publicly available benchmarks.
@@ 4 Experiments
In this section, we conduct extensive experiments to evaluate the effectiveness of our proposed DeLiVR model. We first introduce the experimental setup, including the datasets, evaluation metrics, and implementation details. Subsequently, we compare DeLiVR with several state-of-the-art (SOTA) video deraining methods through quantitative and qualitative analyses on both synthetic and real-world datasets. Furthermore, we conduct a thorough ablation study to validate the contribution of each key component in our model. Finally, we evaluate the impact of deraining on downstream high-level vision tasks—object detection and semantic segmentation—to demonstrate the practical value of our method.
@@ 4.1 Experimental Setup
Our DeLiVR model is implemented using the PyTorch framework and trained on 8 NVIDIA 3090 GPUs. We use the AdamW optimizer with an initial learning rate of $2\times 10^{-4}$. The learning rate is decayed over 5000 epochs using a cosine annealing schedule. We set the batch size to 64 and enable automatic mixed precision (AMP) for training efficiency. For each input, we randomly sample a window of $T=5$ consecutive frames from the video sequences and apply a random crop to a spatial resolution of $256\times 256$. Our model is trained with an L1 reconstruction loss, aiming to restore the clean center frame from the rainy input sequence.
@@ 4.2.1 Quantitative Analysis
As shown in Table 1, DeLiVR consistently achieves the best overall performance across synthetic and real rainy benchmarks. On NTURain, VDMamba (CVPR 25) reports slightly higher PSNR (36.29 vs. 34.06) and SSIM (0.973 vs. 0.952), but at the cost of poorer perceptual quality (LPIPS 0.010 vs. 0.039). More importantly, when evaluated on the real-world WeatherBench dataset, VDMamba suffers a noticeable performance drop (PSNR 23.91, SSIM 0.773), while DeLiVR establishes new state-of-the-art results (PSNR 26.56, SSIM 0.781). This contrast highlights that our Lie-group differential bias design not only provides competitive results on synthetic benchmarks but also yields superior generalization to challenging real rainy scenarios, demonstrating the practical robustness of our approach.
Table 1: Quantitative comparison on synthetic benchmarks and the real-world WeatherBench dataset, including the LPIPS perceptual metric. The WeatherBench benchmark is included to specifically evaluate the model’s generalization ability to authentic, real-world adverse weather conditions. For PSNR/SSIM, higher is better ($\uparrow$). For LPIPS, lower is better ($\downarrow$). Best results are in bold, and second-best are underlined.
@@ 4.2.2 Qualitative Analysis
Figure 3: Qualitative comparison with state-of-the-art methods on four benchmarks. From top to bottom, the rows show results on NTU, Rain-Syn-Light (Light), Rain-Syn-Complex (Complex), and the real-world WeatherBench dataset. Compared to other methods, our model more effectively removes severe rain streaks and color casts (e.g., the red artifacts in the first row), while better preserving fine background textures and natural colors (e.g., rows 2 and 3). Visually, our results are the closest to the Ground Truth (GT).
@@ A.2 Core Pseudocode of DeLiVR
A.3 Use of Large Language Models

### v2

@@ Abstract
Videos captured in the wild often suffer from rain streaks, blur, and noise. In addition, even slight changes in camera pose can amplify cross-frame mismatches and temporal artifacts. Existing methods rely on optical flow or heuristic alignment, which are computationally expensive and less robust. To address these challenges, Lie groups provide a principled way to represent continuous geometric transformations, making them well-suited for enforcing spatial and temporal consistency in video modeling. Building on this insight, we propose DeLiVR, an efficient video deraining method that injects spatiotemporal Lie-group differential biases directly into attention scores of the network. Specifically, the method introduces two complementary components. First, a rotation-bounded Lie relative bias predicts the in-plane angle of each frame using a compact prediction module, where normalized coordinates are rotated and compared with base coordinates to achieve geometry-consistent alignment before feature aggregation. Second, a differential group displacement computes angular differences between adjacent frames to estimate a velocity. This bias computation combines temporal decay and attention masks to focus on inter-frame relationships while precisely matching the direction of rain streaks. Extensive experimental results demonstrate the effectiveness of our method on publicly available benchmarks. The code is publicly available at https://github.com/Shuning0312/ICLR-DeLiVR.
@@ 4.1 Experimental Setup
Our DeLiVR model is implemented using the PyTorch framework and trained on 8 NVIDIA 3090 GPUs. We use the AdamW optimizer with an initial learning rate of $2\times 10^{-4}$. The learning rate is decayed over 5000 epochs using a cosine annealing schedule. We set the batch size to 64. For each input, we randomly sample a window of $T=5$ consecutive frames from the video sequences. Our model is trained with an L1 reconstruction loss, aiming to restore the clean center frame from the rainy input sequence.
@@ 4.2.1 Quantitative Analysis
As shown in Table 1, DeLiVR consistently achieves the best overall performance across synthetic and real rainy benchmarks. On NTURain, our method ranks second only to VDMamba (CVPR 2025). More importantly, when evaluated on the real-world WeatherBench dataset, VDMamba suffers a noticeable performance drop (PSNR 23.91, SSIM 0.773), while DeLiVR establishes new state-of-the-art results (PSNR 26.56, SSIM 0.781). This contrast highlights that our Lie-group differential bias design not only provides competitive results on synthetic benchmarks but also yields superior generalization to challenging real rainy scenarios, demonstrating the practical robustness of our approach.
Table 1: Quantitative comparison on synthetic benchmarks and the real-world WeatherBench dataset. The WeatherBench benchmark is included to specifically evaluate the model’s generalization ability to authentic, real-world adverse weather conditions. For PSNR/SSIM, higher is better ($\uparrow$). For LPIPS, lower is better ($\downarrow$). Best results are in bold, and second-best are underlined.
@@ 4.2.2 Qualitative Analysis
Figure 3: Qualitative comparison with state-of-the-art methods on four benchmarks. From top to bottom, the rows show results on NTU, Rain-Syn-Light, Rain-Syn-Complex, and the real-world WeatherBench dataset. Compared to other methods, our model more effectively removes severe rain streaks and color casts, while better preserving fine background textures and natural colors.
@@ 4.3 Ablation Study
Figure 4: Visual ablation of different bias components on the NTURain dataset. From left to right: baseline without Lie bias, model with spatial bias only (Bspace), model with temporal bias only (Btime), and the full DeLiVR with both spatial and temporal Lie-group differential biases.
@@
Rotation–Perturbation Study. To directly validate that the proposed Lie-group differential bias improves cross-frame alignment, we conduct a controlled experiment where small in-plane rotations are injected into several frames of the input sequence. This perturbation produces synthetic misalignment while keeping scene content unchanged. As shown in Fig. 5, the rotation-aware model exhibits clearer temporal stability. The attention maps further reveal that the rotation-enhanced model captures richer directional dependencies, displaying higher entropy and more concentrated high-valued regions. The attention-difference visualization highlights that the Lie-group bias preferentially strengthens correspondences around motion-sensitive structures, thereby reducing flicker and ghosting. These results provide explicit interpretability evidence that our method addresses cross-frame misalignment beyond improving overall metrics.
Controlled Comparison with Optical Flow Biases. To provide a comparison between our Lie-group rotation module and optical-flow-based alternatives, we conducted a strictly controlled experiment where all variables except the bias representation were held constant. Specifically, we utilized a 12-layer Transformer backbone and maintained an identical training protocol. For the optical-flow baseline, RAFT flow fields were processed through a small learnable MLP to predict bias maps, mirroring the structure of our rotation module. This design ensures the only differing factor is the choice of bias representation. Evaluations on the NTU-Rain dataset show our method improves PSNR by +2.43 dB, demonstrating that modeling rain-streak orientation on the $SO(2)$ manifold provides a stronger and more stable geometric inductive bias compared with unconstrained optical-flow fields. Visualizations in in Appendix (Fig. 9) and real-world results in Appendix (Fig. 10) further confirm that our proposed prior excels on benchmarks and transfers robustly to naturally occurring rain conditions.
Figure 5: Attention comparison between the baseline and the rotation-enhanced model. The top row shows full attention matrices for both models under the same input sequence. The second row visualizes spatial attention maps extracted from representative query positions. The right column reports aggregated attention statistics, including average entropy and maximum attention values. The bottom row presents the distribution of attention weights and attention entropy across all heads and layers.
@@ 4.4 Discussion
Comparison with Optical-Flow-Based Guidance. We evaluate two representative flow-guided recurrent video restoration models on the WeatherBench dataset: Frame-Consistent Recurrent Video Deraining (Yang et al., 2019) (PSNR 23.33 / SSIM 0.691 / LPIPS 0.455) and High-Resolution Optical Flow and Frame-Recurrent Network (Fang & Zhan, 2022) (PSNR 24.10 / SSIM 0.716 / LPIPS 0.402). Under the same evaluation setting on an A100 GPU, our rotation-prediction-based model consistently surpasses both flow-based baselines.
@@ A.2 Core Pseudocode of DeLiVR
A.3 Generalization to Deblurring and Dehazing
Although our method is introduced primarily for video deraining, the core idea—injecting a Lie-group differential bias to stabilize cross-frame alignment—is degradation-agnostic. To further verify this, we additionally evaluate the model on two representative degradation types: deblurring (trained on the GoPro dataset  (Nah et al., 2017)) and dehazing (trained on the RESIDE datasetLi et al. (2019)). As shown in Fig. 7, our method consistently improves geometric stability, reduces temporal flickering, and preserves fine structures across all three tasks (deraining, deblurring, dehazing). These results demonstrate that the proposed geometric prior captures fundamental motion characteristics rather than degradation-specific cues, thereby enabling natural transfer to a broader range of video restoration scenarios.
Figure 7: Generalization to other degradations. Our method consistently produces cleaner structures and fewer temporal flickers.
A.4 Real-World Downstream Evaluation
To further validate the practical utility of our method, we test an off-the-shelf object detector on a real captured rainy image. As shown in Fig. 8, deraining enables the detector to identify an additional pedestrian missed in the rainy input, demonstrating improved reliability for downstream perception.
Figure 8: Real-world downstream evaluation on object detection. We apply an off-the-shelf detector to a real captured rainy frame (left) and to the derained result produced by our method (right).
Figure 9: Visual comparison on the NTU-Rain. From left to right: rainy input, flow-based bias baseline, our Lie Rotation model, and ground-truth.
Figure 10: Real-world rain removal. From left to right: the original rainy input, the flow-based bias baseline, and our Lie Rotation model.
A.5 Large Visualization of Figure 3
To provide a clearer view, we include the enlarged version of Figure 3 in Appendix.
Figure 11: Enlarged visualization corresponding to Figure 3.
A.6 Use of Large Language Models